# sangkuriang: A pseudo-spectral Python library for Korteweg-de Vries soliton simulation

### v2

@@ 3 Results and Discussion
Case NN $\Delta x$ [m] $T_{\text{final}}$ [s] Steps CPU time [s] Steps/s
1 (Single soliton) 512 0.117 50 208,202 152.0 1370
2 (Two equal) 512 0.157 60 98,612 75.7 1303
3 (Collision) 512 0.157 70 115,658 87.6 1320
4 (Three solitons) 1024 0.098 80 551,858 534.4 1033
The throughput, measured in time steps per second, remained relatively consistent across Cases 1–3 at approximately 1300–1370 steps/s, decreasing to roughly 1030 steps/s for Case 4. This reduction likely reflects the increased memory bandwidth demands associated with the larger grid and the superlinear scaling of FFT operations, which nominally follow $\mathcal{O}(N\log N)$ complexity [30]. Nevertheless, all simulations completed within times amenable to interactive research workflows.
The preservation of the three KdV invariants—mass $M$, momentum $P$, and energy $E$ as defined in equations (45)–(47)—provides a stringent test of numerical accuracy. Figure 2 displays the temporal evolution of relative conservation errors for all test cases.
@@
The single-soliton case exhibited the smallest conservation errors, with maximum relative deviations of $7.64\times 10^{-5}$ for mass, $2.98\times 10^{-7}$ for momentum, and $1.73\times 10^{-7}$ for energy. These values are consistent with the specified integration tolerances and confirm that the pseudo-spectral spatial discretization introduces negligible additional error for smooth, non-interacting solutions. The root-mean-square (RMS) errors were approximately half the maximum values ($3.55\times 10^{-5}$, $1.30\times 10^{-7}$, and $8.28\times 10^{-8}$, respectively), indicating that the deviations represent bounded oscillations rather than secular drift. Linear regression of the conservation time series yielded drift rates on the order of $10^{-9}$ s-1 for momentum and energy, suggesting that the dominant error contribution arises from the non-symplectic character of the DOP853 integrator rather than from systematic bias [22].
@@
Multi-soliton configurations exhibited moderately larger conservation errors, as anticipated from the increased solution complexity during interaction events. The two equal-amplitude soliton case (Case 2) showed maximum errors of $1.57\times 10^{-4}$ (mass), $1.90\times 10^{-6}$ (momentum), and $1.45\times 10^{-6}$ (energy). The overtaking collision (Case 3) produced the largest mass error at $3.33\times 10^{-4}$, with momentum and energy errors of $6.49\times 10^{-6}$ and $2.60\times 10^{-6}$, respectively. Figure 2(a)–(c) reveals that the conservation deviations exhibit transient excursions correlated with collision events, particularly evident in the three-soliton case where multiple interactions occur. Despite the increased complexity of Case 4, the maximum errors ($2.56\times 10^{-4}$, $8.86\times 10^{-6}$, $6.20\times 10^{-6}$) remained comparable to or smaller than those of Case 3, likely reflecting the benefits of enhanced spatial resolution and tighter tolerances employed for this configuration.
@@
The bar chart in Figure 2(d) presents a comparative summary of maximum relative errors across all cases. A consistent hierarchy emerges: mass conservation errors exceed momentum errors by approximately two orders of magnitude, which in turn exceed energy errors. This ordering may reflect the different sensitivities of each integral to high-wavenumber components of the solution, with the energy functional (32) involving spatial derivatives that weight the well-resolved low-wavenumber content more heavily. Across all test cases, the conservation errors remain well below the $10^{-2}$ threshold commonly adopted as acceptable for practical simulations of Hamiltonian systems [36], suggesting that the sangkuriang implementation maintains adequate fidelity to the underlying physics.
@@
Figure 3 presents three-dimensional visualizations of the wave field $u(x,t)$ for each test case, rendered with consistent axis limits ($x\in[-50,50]$ m, $t\in[0,80]$ s, $u\in[0,9]$ m) to facilitate direct comparison.
@@
Panel (a) illustrates the single-soliton case, where the initial $\text{sech}^{2}$ profile propagates from left to right while maintaining its characteristic shape. The wave amplitude increases slightly from the initial value of approximately 4.0 m to a final peak of 5.2 m, with a global maximum of 5.21 m observed during the simulation. This amplitude evolution reflects the adjustment of the initial condition—which approximates but does not precisely match the exact soliton solution (33)—toward a true soliton state through radiation of small-amplitude dispersive waves [6]. The peak position advanced by 17.8 m over the 50 s integration interval, corresponding to an average velocity of 0.357 m/s.
@@
The two equal-amplitude soliton configuration, shown in panel (b), exhibits parallel propagation of two $A=3.0$ m solitons initially separated by 30 m. Since the KdV soliton velocity is proportional to amplitude according to equation (34), these waves maintain constant separation throughout the simulation. The surface plot reveals two distinct ridges traversing the space-time domain without intersection, confirming the expected kinematic behavior. Minor amplitude variations (final peak 3.66 m versus initial 3.0 m) again indicate adjustment toward exact soliton profiles.
@@
Panel (c) displays the overtaking collision scenario, which constitutes the most stringent test of the solver’s ability to capture nonlinear interaction dynamics. The taller soliton ($A_{1}=6.0$ m), propagating faster than the shorter wave ($A_{2}=2.0$ m), overtakes and passes through it near $t\approx 35$ s and $x\approx 0$ m. The characteristic “X”-shaped intersection pattern in the space-time surface demonstrates the elastic scattering property of KdV solitons: both waves emerge from the collision with their identities intact, differing only by phase shifts [2]. The global maximum amplitude of 7.55 m, exceeding the linear superposition of the two initial amplitudes, reflects the transient nonlinear amplification during the collision phase.
@@
The three-soliton interaction in panel (d) presents the most complex dynamics, with multiple collision events occurring as the fastest soliton ($A_{1}=7.0$ m) successively overtakes the intermediate ($A_{2}=4.0$ m) and slowest ($A_{3}=2.5$ m) waves. The space-time surface exhibits a sequence of intersection patterns, with a global maximum of 8.26 m observed during the highest-amplitude collision. Despite this complexity, the three distinct soliton tracks remain identifiable throughout the simulation, consistent with the complete integrability of the KdV equation [4].
@@
Panel (c) synthesizes the velocity–amplitude relationship across all tracked solitons. The data points cluster along the theoretical line $v=\varepsilon A/3$, providing quantitative confirmation that the computed solutions satisfy the fundamental KdV dispersion relation. The agreement is particularly notable given that the initial conditions employ approximate sech2 profiles rather than exact soliton solutions, demonstrating the robustness of the soliton state as an attractor in the KdV phase space [3].
Panel (d) presents a sequence of wave profiles during the collision event, vertically offset for visualization. The temporal progression illustrates the collision mechanism: the faster soliton approaches from behind, the two waves temporarily merge into a single enhanced peak, and subsequently separate with their original amplitudes restored. This elastic scattering behavior, wherein solitons interact nonlinearly yet emerge unchanged, constitutes the defining characteristic that distinguishes solitons from ordinary wave packets [2].
@@
Panel (a) displays the initial power spectral density $P(k)$ on logarithmic axes. All cases exhibit similar spectral structure: a plateau at low wavenumbers ($k<0.1$ rad/m) transitioning to exponential decay at higher wavenumbers, consistent with the smooth sech2 spatial profiles [18]. The three-soliton case (dotted red) shows the highest total power, reflecting the larger combined amplitude, while the two equal-amplitude case (dashed orange) exhibits oscillatory modulation arising from the spatial interference pattern of the two separated solitons. The spectral decay extends over approximately 15 orders of magnitude before reaching the numerical noise floor near $10^{-16}$, confirming that aliasing errors remain negligible for the employed spatial resolutions.
The normalized spectral entropy $S_{k}$, plotted in panel (b), quantifies the breadth of the spectral energy distribution according to equation (63). Values near unity indicate uniform spectral content (approaching white noise), while values near zero correspond to energy concentration at a single wavenumber (monochromatic). The soliton solutions exhibit intermediate values in the range 0.460.460.590.59, reflecting their localized but smooth spatial structure. The single-soliton case shows relatively stable entropy ($0.528\pm 0.021$), while the collision case displays larger fluctuations ($0.567\pm 0.024$) with transient increases during the interaction event near $t=35$ s. These variations suggest temporary spectral broadening as the wave profiles distort during collision, followed by return to the pre-collision distribution as the solitons separate.
@@
Panel (d) presents the Fisher information $F$, which measures the sharpness or localization of the spatial amplitude distribution according to equation (67). Higher values indicate more localized, sharper-peaked structures. The three-soliton case exhibits the highest mean Fisher information ($0.960\pm 0.148$ m-2), reflecting the presence of multiple narrow peaks, while the two equal-amplitude case shows the lowest ($0.641\pm 0.119$ m-2) due to the broader, lower-amplitude solitons. Notably, the collision case displays pronounced spikes in Fisher information (reaching 1.8 m-2) during the interaction event, reflecting the transient steepening of gradients as the soliton profiles merge. These transient features subsequently relax as the waves separate.
@@
The results presented above demonstrate that the sangkuriang solver successfully reproduces the essential features of KdV soliton dynamics while maintaining numerical accuracy appropriate for research applications. Several aspects merit further discussion.
@@
The pseudo-spectral spatial discretization combined with eighth-order adaptive time integration achieved conservation errors ranging from $\mathcal{O}(10^{-7})$ to $\mathcal{O}(10^{-4})$ across the test cases. These values compare favorably with those reported in the literature for similar methods applied to nonlinear dispersive equations [17, 20]. The observation that momentum and energy are conserved more accurately than mass warrants consideration. The mass functional (30) involves only the solution field itself, making it sensitive to any drift in the solution mean. In contrast, the momentum functional (31) depends on $u^{2}$, which suppresses the contribution of small-amplitude dispersive radiation relative to the dominant soliton peaks. The energy functional (32), involving spatial derivatives, further weights the smooth soliton cores over high-wavenumber numerical artifacts. This hierarchy suggests that the mass error may serve as a conservative upper bound on overall numerical accuracy.
@@
The absence of systematic drift in the conservation time series indicates that the DOP853 integrator, while not symplectic, introduces primarily oscillatory rather than secular errors for the integration times considered. For significantly longer simulations, the accumulation of these oscillatory errors could potentially become problematic, and structure-preserving integrators [36, 44] might offer advantages despite their typically higher computational cost per step.
@@
The close agreement between measured and theoretical soliton velocities ($5\%$ discrepancy) confirms that the computed solutions satisfy the fundamental amplitude–velocity relationship (34). The small systematic excess in measured velocities relative to predictions based on mean amplitude may arise from several factors. First, the soliton amplitudes evolve during the initial adjustment phase as the approximate initial conditions relax toward exact soliton profiles, and velocity measurements based on trajectory slopes integrate over this transient. Second, the theoretical velocity assumes an isolated soliton in an infinite domain, whereas the periodic boundary conditions and finite domain introduce corrections that may slightly modify the effective dispersion relation [18].
@@
The high linearity of all tracked trajectories ($R^{2}>0.99$) provides strong evidence that the computed waves behave as true solitons rather than dispersive wave packets that would spread and decelerate over time. This validation is particularly meaningful for the post-collision segments, where any deviation from integrability would manifest as amplitude-dependent velocity changes.
@@
The spectral and information-theoretic analyses reveal consistent patterns that reflect the integrable structure of the KdV equation. The bounded variations of spectral entropy, statistical complexity, and Fisher information—with transient excursions during collisions that subsequently relax—indicate that the wave field returns to its pre-collision information content. This behavior contrasts with what would be expected for non-integrable systems, where collision events could transfer energy to radiation modes, leading to secular changes in these measures [6].
@@
The Fisher information proved particularly sensitive to collision dynamics, with pronounced spikes corresponding to the transient steepening of gradients during wave merging. This sensitivity suggests that Fisher information may serve as a useful diagnostic for detecting the onset and completion of soliton interactions in more complex scenarios.
@@
These results suggest that RQA could serve as a model-independent diagnostic for assessing the integrability of computed solutions in situations where analytical results are unavailable. Significant deviations from high determinism might indicate either numerical breakdown of integrability due to insufficient resolution, or genuine non-integrable dynamics in perturbed KdV-type equations.
@@
The demonstration that research-grade KdV simulations can be performed in minutes on a laptop-class workstation represents a practical advantage for exploratory studies. The Numba-accelerated implementation achieved throughputs of approximately 1000–1400 time steps per second, suggesting that the solver is not severely memory-bandwidth limited for the grid sizes considered. The observed scaling from $N=512$ to $N=1024$ (throughput reduction from $\sim$1300 to $\sim$1030 steps/s) is consistent with the $\mathcal{O}(N\log N)$ FFT complexity combined with increased cache pressure.
@@
For applications requiring higher resolution or longer integration times, several optimization strategies could be considered: vectorized evaluation of multiple initial conditions in parallel, GPU acceleration of FFT operations via CuPy [45], or time-parallel algorithms for ensemble simulations. The modular architecture of sangkuriang is designed to facilitate such extensions while maintaining compatibility with the existing analysis framework.
@@
Several limitations of the present study should be acknowledged. First, the test cases employed approximate sech2 initial conditions rather than exact $N$-soliton solutions obtainable from inverse scattering theory [5]; comparison with exact solutions would provide a more stringent validation. Second, the periodic boundary conditions preclude investigation of soliton interactions with boundaries or non-periodic wave backgrounds. Third, the standard KdV equation considered here represents only the leading-order balance between nonlinearity and dispersion; extensions to higher-order equations (e.g., the modified KdV or Kawahara equations [46]) would broaden the solver’s applicability to different physical regimes.
@@
Future development of sangkuriang could address these limitations while maintaining the design philosophy of accessibility and reproducibility. The integration of inverse scattering routines would enable exact initial conditions for validation studies. Variable-coefficient extensions would permit modeling of wave propagation over varying topography. Perhaps most significantly, the pseudo-spectral framework could be extended to two-dimensional generalizations such as the Kadomtsev–Petviashvili equation [6], enabling investigation of oblique soliton interactions and transverse instabilities.

### v3

@@ 3 Results
Case NN $\Delta x$ [m] $T_{\text{final}}$ [s] Steps CPU Time [s]
1 (Single soliton) 512 0.117 50 208,202 152.0
2 (Two equal) 512 0.157 60 98,612 75.7
3 (Collision) 512 0.157 70 115,658 87.6
4 (Three solitons) 1024 0.098 80 551,858 534.4
The throughput, measured in time steps per second, remained relatively consistent across Cases 1–3 at approximately 1300–1370 steps/s, decreasing to roughly 1030 steps/s for Case 4. This reduction likely reflects the increased memory bandwidth demands associated with the larger grid and the superlinear scaling of FFT operations, which nominally follow $\mathcal{O}(N\log N)$ complexity [30]. Nevertheless, all simulations completed within times amenable to interactive research workflows, enabling rapid parameter exploration in idealized ocean soliton studies.
The preservation of the three KdV invariants—mass $M$, momentum $P$, and energy $E$ as defined in equations (45)–(47)—provides a stringent test of numerical accuracy. For ocean wave applications, the faithful conservation of these quantities ensures that the solver does not introduce spurious wave energy sources or sinks that could contaminate the simulated dynamics [33]. Figure 2 displays the temporal evolution of relative conservation errors for all test cases.
@@
The single-soliton case exhibited the smallest conservation errors, with maximum relative deviations of $7.64\times 10^{-5}$ for mass, $2.98\times 10^{-7}$ for momentum, and $1.73\times 10^{-7}$ for energy. These values are consistent with the specified integration tolerances and confirm that the pseudo-spectral spatial discretization introduces negligible additional error for smooth, non-interacting solutions representative of an isolated oceanic solitary wave. The root-mean-square (RMS) errors were approximately half the maximum values ($3.55\times 10^{-5}$, $1.30\times 10^{-7}$, and $8.28\times 10^{-8}$, respectively), indicating that the deviations represent bounded oscillations rather than secular drift. Linear regression of the conservation time series yielded drift rates on the order of $10^{-9}$ s-1 for momentum and energy, suggesting that the dominant error contribution arises from the non-symplectic character of the DOP853 integrator rather than from systematic bias [22].
@@
Multi-soliton configurations exhibited moderately larger conservation errors, as anticipated from the increased solution complexity during interaction events. The two equal-amplitude soliton case (Case 2) showed maximum errors of $1.57\times 10^{-4}$ (mass), $1.90\times 10^{-6}$ (momentum), and $1.45\times 10^{-6}$ (energy). The overtaking collision (Case 3) produced the largest mass error at $3.33\times 10^{-4}$, with momentum and energy errors of $6.49\times 10^{-6}$ and $2.60\times 10^{-6}$, respectively. Figure 2a–c reveals that the conservation deviations exhibit transient excursions correlated with collision events, particularly evident in the three-soliton case where multiple interactions occur. Such transient increases in conservation error during wave–wave interaction are expected physically, as the nonlinear coupling intensifies when soliton profiles overlap [6]. Despite the increased complexity of Case 4, the maximum errors ($2.56\times 10^{-4}$, $8.86\times 10^{-6}$, $6.20\times 10^{-6}$) remained comparable to or smaller than those of Case 3, likely reflecting the benefits of enhanced spatial resolution and tighter tolerances employed for this configuration.
@@
The bar chart in Figure 2d presents a comparative summary of maximum relative errors across all cases. A consistent hierarchy emerges: mass conservation errors exceed momentum errors by approximately two orders of magnitude, which in turn exceed energy errors. This ordering may reflect the different sensitivities of each integral to high-wavenumber components of the solution, with the energy functional (32) involving spatial derivatives that weight the well-resolved low-wavenumber content more heavily. Across all test cases, the conservation errors remain well below the $10^{-2}$ threshold commonly adopted as acceptable for practical simulations of Hamiltonian systems [33], suggesting that the sangkuriang implementation maintains adequate fidelity to the underlying physics governing idealized ocean soliton propagation.
@@
Figure 3 presents three-dimensional visualizations of the wave field $u(x,t)$ for each test case, rendered with consistent axis limits ($x\in[-50,50]$ m, $t\in[0,80]$ s, $u\in[0,9]$ m) to facilitate direct comparison of the idealized oceanic soliton dynamics.
@@
Panel (a) illustrates the single-soliton case, where the initial $\text{sech}^{2}$ profile propagates from left to right while maintaining its characteristic shape—the hallmark of a solitary wave in a dispersive medium. The wave amplitude increases slightly from the initial value of approximately 4.0 m to a final peak of 5.2 m, with a global maximum of 5.21 m observed during the simulation. This amplitude evolution reflects the adjustment of the initial condition—which approximates but does not precisely match the exact soliton solution (33)—toward a true soliton state through radiation of small-amplitude dispersive waves [6]. Such adjustment dynamics are analogous to the fission process observed when a tidal bore or internal tide evolves into a rank-ordered train of solitary waves on a continental shelf [46, 49]. The peak position advanced by 17.8 m over the 50 s integration interval, corresponding to an average velocity of 0.357 m/s.
@@
The two equal-amplitude soliton configuration, shown in panel (b), exhibits parallel propagation of two $A=3.0$ m solitons initially separated by 30 m. Since the KdV soliton velocity is proportional to amplitude according to equation (34), these waves maintain constant separation throughout the simulation, akin to the approximately equal-amplitude internal waves sometimes observed in the trailing portion of oceanic soliton packets [48]. The surface plot reveals two distinct ridges traversing the space-time domain without intersection, confirming the expected kinematic behavior. Minor amplitude variations (final peak 3.66 m versus initial 3.0 m) again indicate adjustment toward exact soliton profiles.
@@
Panel (c) displays the overtaking collision scenario, which constitutes the most stringent test of the solver’s ability to capture nonlinear interaction dynamics. The taller soliton ($A_{1}=6.0$ m), propagating faster than the shorter wave ($A_{2}=2.0$ m), overtakes and passes through it near $t\approx 35$ s and $x\approx 0$ m. The characteristic “X”-shaped intersection pattern in the space-time surface demonstrates the elastic scattering property of KdV solitons: both waves emerge from the collision with their identities intact, differing only by phase shifts [3]. This elastic interaction is the idealized analog of the overtaking collisions documented in field observations of internal solitary waves, where larger waves catch and pass smaller ones with remarkably little energy exchange [47, 51]. The global maximum amplitude of 7.55 m, exceeding the linear superposition of the two initial amplitudes, reflects the transient nonlinear amplification during the collision phase.
@@
The three-soliton interaction in panel (d) presents the most complex dynamics, with multiple collision events occurring as the fastest soliton ($A_{1}=7.0$ m) successively overtakes the intermediate ($A_{2}=4.0$ m) and slowest ($A_{3}=2.5$ m) waves. The space-time surface exhibits a sequence of intersection patterns, with a global maximum of 8.26 m observed during the highest-amplitude collision. Despite this complexity, the three distinct soliton tracks remain identifiable throughout the simulation, consistent with the complete integrability of the KdV equation [4]. This rank-ordered configuration, with the largest and fastest wave leading, mirrors the structure of oceanic internal soliton packets routinely observed in satellite synthetic aperture radar imagery [46, 49].
@@
Panel (c) synthesizes the velocity–amplitude relationship across all tracked solitons. The data points cluster along the theoretical line $v=\varepsilon A/3$, providing quantitative confirmation that the computed solutions satisfy the fundamental KdV dispersion relation. The agreement is particularly notable given that the initial conditions employ approximate sech2 profiles rather than exact soliton solutions, demonstrating the robustness of the soliton state as an attractor in the KdV phase space [1]. This amplitude–velocity scaling is the same relationship that governs the rank ordering and separation of internal solitary waves observed in the ocean [46, 48].
Panel (d) presents a sequence of wave profiles during the collision event, vertically offset for visualization. The temporal progression illustrates the collision mechanism: the faster soliton approaches from behind, the two waves temporarily merge into a single enhanced peak, and subsequently separate with their original amplitudes restored. This elastic scattering behavior, wherein solitons interact nonlinearly yet emerge unchanged, constitutes the defining characteristic that distinguishes solitons from ordinary dispersive wave packets [3].
@@
Panel (a) displays the initial power spectral density $P(k)$ on logarithmic axes. All cases exhibit similar spectral structure: a plateau at low wavenumbers ($k<0.1$ rad/m) transitioning to exponential decay at higher wavenumbers, consistent with the smooth sech2 spatial profiles characteristic of KdV solitons [18]. The three-soliton case (dotted red) shows the highest total power, reflecting the larger combined amplitude, while the two equal-amplitude case (dashed orange) exhibits oscillatory modulation arising from the spatial interference pattern of the two separated solitons. The spectral decay extends over approximately 15 orders of magnitude before reaching the numerical noise floor near $10^{-16}$, confirming that aliasing errors remain negligible for the employed spatial resolutions.
The normalized spectral entropy $S_{k}$, plotted in panel (b), quantifies the breadth of the spectral energy distribution according to equation (63). Values near unity indicate uniform spectral content (approaching white noise), while values near zero correspond to energy concentration at a single wavenumber (monochromatic). The soliton solutions exhibit intermediate values in the range 0.460.460.590.59, reflecting their localized but smooth spatial structure. The single-soliton case shows relatively stable entropy ($0.528\pm 0.021$), while the collision case displays larger fluctuations ($0.567\pm 0.024$) with transient increases during the interaction event near $t=35$ s. These variations suggest temporary spectral broadening as the wave profiles distort during collision, followed by return to the pre-collision distribution as the solitons separate—a spectral signature of the elastic scattering process.
@@
Panel (d) presents the Fisher information $F$, which measures the sharpness or localization of the spatial amplitude distribution according to equation (67). Higher values indicate more localized, sharper-peaked structures. The three-soliton case exhibits the highest mean Fisher information ($0.960\pm 0.148$ m-2), reflecting the presence of multiple narrow peaks, while the two equal-amplitude case shows the lowest ($0.641\pm 0.119$ m-2) due to the broader, lower-amplitude solitons. Notably, the collision case displays pronounced spikes in Fisher information (reaching 1.8 m-2) during the interaction event, reflecting the transient steepening of gradients as the soliton profiles merge. These transient features subsequently relax as the waves separate, consistent with the elastic nature of KdV soliton collisions.
@@ 4 Discussion
The results presented above demonstrate that the sangkuriang solver successfully reproduces the essential features of KdV soliton dynamics while maintaining numerical accuracy appropriate for research applications in idealized ocean wave modeling. Several aspects merit further discussion in the context of GFD and coastal oceanography.
@@
The pseudo-spectral spatial discretization combined with eighth-order adaptive time integration achieved conservation errors ranging from $\mathcal{O}(10^{-7})$ to $\mathcal{O}(10^{-4})$ across the test cases. These values compare favorably with those reported in the literature for similar methods applied to nonlinear dispersive equations arising in ocean wave theory [17, 20]. The observation that momentum and energy are conserved more accurately than mass warrants consideration. The mass functional (30) involves only the solution field itself, making it sensitive to any drift in the solution mean. In contrast, the momentum functional (31) depends on $u^{2}$, which suppresses the contribution of small-amplitude dispersive radiation relative to the dominant soliton peaks. The energy functional (32), involving spatial derivatives, further weights the smooth soliton cores over high-wavenumber numerical artifacts. This hierarchy suggests that the mass error may serve as a conservative upper bound on overall numerical accuracy. From an oceanographic standpoint, the high fidelity of momentum and energy conservation is particularly reassuring, as these quantities govern the transport of wave energy and the forces exerted by solitary waves on offshore structures and the seabed [48].
@@
The absence of systematic drift in the conservation time series indicates that the DOP853 integrator, while not symplectic, introduces primarily oscillatory rather than secular errors for the integration times considered. For significantly longer simulations—such as those required to model solitary wave propagation across an entire continental shelf or ocean basin—the accumulation of these oscillatory errors could potentially become problematic, and structure-preserving integrators [33, 43] might offer advantages despite their typically higher computational cost per step.
@@
The close agreement between measured and theoretical soliton velocities ($5\%$ discrepancy) confirms that the computed solutions satisfy the fundamental amplitude–velocity relationship (34). This nonlinear dispersion relation is the same scaling that governs the rank ordering and separation of internal solitary waves observed in the ocean, where larger-amplitude waves consistently propagate faster than their smaller counterparts [46, 49]. The small systematic excess in measured velocities relative to predictions based on mean amplitude may arise from several factors. First, the soliton amplitudes evolve during the initial adjustment phase as the approximate initial conditions relax toward exact soliton profiles, and velocity measurements based on trajectory slopes integrate over this transient. Second, the theoretical velocity assumes an isolated soliton in an infinite domain, whereas the periodic boundary conditions and finite domain introduce corrections that may slightly modify the effective dispersion relation [18]. Similar discrepancies between observed and theoretically predicted propagation speeds have been documented in field studies of oceanic internal solitary waves, where background currents, stratification variability, and bathymetric effects introduce additional corrections to the idealized KdV prediction [47, 50].
@@
The high linearity of all tracked trajectories ($R^{2}>0.99$) provides strong evidence that the computed waves behave as true solitons rather than dispersive wave packets that would spread and decelerate over time. This validation is particularly meaningful for the post-collision segments, where any deviation from integrability would manifest as amplitude-dependent velocity changes. In the oceanic context, the persistence of soliton identity through collisions is a key feature that enables the long-range propagation of internal solitary waves across hundreds of kilometers of continental shelf [46, 51].
@@
The spectral and information-theoretic analyses reveal consistent patterns that reflect the integrable structure of the KdV equation. The bounded variations of spectral entropy, statistical complexity, and Fisher information—with transient excursions during collisions that subsequently relax—indicate that the wave field returns to its pre-collision information content. This behavior contrasts with what would be expected for non-integrable systems, where collision events could transfer energy to radiation modes, leading to secular changes in these measures [6]. In realistic ocean settings, perturbations to perfect integrability arise from variable stratification, background shear, and bathymetric gradients [50]; the diagnostic tools developed here could serve to quantify the degree to which such perturbations disrupt the regular structure of the idealized KdV dynamics.
@@
The Fisher information proved particularly sensitive to collision dynamics, with pronounced spikes corresponding to the transient steepening of gradients during wave merging. This sensitivity suggests that Fisher information may serve as a useful diagnostic for detecting the onset and completion of soliton interactions in more complex scenarios, including the multi-wave interaction events observed in energetic oceanic internal wave environments such as the South China Sea [49] and the Strait of Gibraltar.
@@
These results suggest that RQA could serve as a model-independent diagnostic for assessing the integrability of computed solutions in situations where analytical results are unavailable. Significant deviations from high determinism might indicate either numerical breakdown of integrability due to insufficient resolution, or genuine non-integrable dynamics in perturbed KdV-type equations. This capability is potentially valuable for ocean wave modeling, where extended KdV equations with variable coefficients are routinely employed to account for realistic oceanographic conditions [48, 50], and where the degree to which the integrable structure survives such extensions is not always known a priori.
@@
The demonstration that research-grade KdV simulations can be performed in minutes on a laptop-class workstation represents a practical advantage for exploratory studies in physical oceanography and GFD. The Numba-accelerated implementation achieved throughputs of approximately 1000–1400 time steps per second, suggesting that the solver is not severely memory-bandwidth limited for the grid sizes considered. The observed scaling from $N=512$ to $N=1024$ (throughput reduction from $\sim$1300 to $\sim$1030 steps/s) is consistent with the $\mathcal{O}(N\log N)$ FFT complexity combined with increased cache pressure. The accessibility of this computational tool on commodity hardware lowers the barrier for researchers and graduate students in oceanography to conduct idealized numerical experiments on nonlinear dispersive wave dynamics without requiring institutional high-performance computing allocations.
@@
For applications requiring higher resolution or longer integration times—such as simulating solitary wave propagation over shelf-scale distances or conducting ensemble studies across a range of oceanic stratification profiles—several optimization strategies could be considered: vectorized evaluation of multiple initial conditions in parallel, GPU acceleration of FFT operations via CuPy [44], or time-parallel algorithms for ensemble simulations. The modular architecture of sangkuriang is designed to facilitate such extensions while maintaining compatibility with the existing analysis framework.
@@
Several limitations of the present study should be acknowledged. First, the test cases employed approximate sech2 initial conditions rather than exact $N$-soliton solutions obtainable from inverse scattering theory [5]; comparison with exact solutions would provide a more stringent validation. Second, the periodic boundary conditions preclude investigation of soliton interactions with boundaries or non-periodic wave backgrounds, which are relevant to realistic coastal settings where waves encounter shorelines, headlands, or river plumes. Third, the standard KdV equation considered here represents only the leading-order balance between nonlinearity and dispersion; extensions to higher-order equations (e.g., the modified KdV or Kawahara equations [45]) and variable-coefficient formulations that account for slowly varying bathymetry and stratification [50] would broaden the solver’s applicability to different oceanic regimes. Fourth, the present study considers only surface wave analogs; the extension to internal wave dynamics in continuously stratified fluids requires the incorporation of modal structure and density-dependent coefficients [46, 48].
@@
Future development of sangkuriang could address these limitations while maintaining the design philosophy of accessibility and reproducibility. The integration of inverse scattering routines would enable exact initial conditions for validation studies. Variable-coefficient extensions would permit modeling of wave propagation over varying bathymetry, a critical capability for realistic coastal ocean applications [50, 51]. The incorporation of rotational effects through the Ostrovsky equation would extend applicability to internal waves at scales where the Earth’s rotation becomes significant [47]. Perhaps most significantly, the pseudo-spectral framework could be extended to two-dimensional generalizations such as the Kadomtsev–Petviashvili equation [6], enabling investigation of oblique soliton interactions and transverse instabilities relevant to the three-dimensional propagation patterns of oceanic internal waves observed in satellite imagery [49].